OpenStudy (anonymous):

Find the two straight lines that are perpendicular to y=0.25x and tangent to the curve f(x)=1/x?

OpenStudy (anonymous):

6x^2 - 14x + 8 = 0 2(3x^2 - 7x + 4) = 0 (3x - 4)(x - 1) so the 2 x coordinates are 4/3 and 1

OpenStudy (noelgreco):

All lines perpendicular to y=.25x have the same slope, that is -1/m